OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of eighty-one and five-sevenths inches. What is its diameter? Use twenty-two sevenths for pi. twelve inches eighteen inches twenty-one inches twenty-six inches

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

Ok the area for circumference is C=2pi r which is the same as C=pi * d d=diameter 81 and 5/7 =572/7 572/7=22/7d multiply both sides by 7/22 d=26 in